The S&OP team at Kansas Furniture, has received estimates of demand requirements as shown in the table. Assuming one-time stockout costs for lost sale per unit, inventory carrying costs of $30 per unit per month, and zero beginning and ending inventory, evaluate these two plans on an incremental cost basis: Plan A: Produce at a steady rate (equal to minimum requirements) of 1,100 units per month and subcontract additional units at a $60 per unit premium cost. Subcontracting capacity is limited to 800 units per month. (Enter all responses as whole numbers). Ending Subcontract Month Demand Production Inventory(Units 1 July 1300 ,100 2 August 3 September 100 ,100 4 October 5 Novembe 900 100 6 December 1900.100 1150 1,100 1600 1,100 The total cost, excluding normal time labor costs, for Plan A (Enter your response as a whole number)
lt 35 pe is workers additional of hiring The cost in June. Plan B: Vary the workforce to produce the prior month's demand. The fim produced 1,300 units The cost of layoffs is $60 per unit cut back. (Enter all responses as whole numbers) e, going from production of 1.300 in July to 1300 in August requires a layoff (and related i costs) of 0 units in August) Hire Layoff Ending Stockouts Month Demand Production Units) (Units)Inventory(Units) 1300 1150 1July 2 August 3 September 1100 4 October 5 November 00 6 December 90 1600 he total hiring cost-(Enter your response as a whole number.) he total layoff cost-$(Enter your response as a whole number)
VUhtUnits) 2 August 1150 3 September 100 1600 5 November 900 6 December 900 4 October The totai hiring cost (Enter your response as a whole number.) The total layoff cost (Enter your response as a whole number.) The total inventory carrying cost-s(Enter your response as a whole number) The total stockout cost (Enter your response as a whole number) The total cost, excluding normal time labor costs, for Plan B s(Enter your response as a whole number)
Answer #1

Plan A

Month

Demand

Production

Ending Inventory

Subcontract(units)

July

1300

1100

0

200

August

1150

1100

0

50

September

1100

1100

0

0

October

1600

1100

0

500

November

1900

1100

0

800

December

1900

1100

0

800

Total

2350

Total cost, excluding normal time labor costs for plan A = Inventory carrying cost + Subcontracting cost

= 0 + 60*2350

= 141000

Plan B

Month

Demand

Production

Hire (Units)

Layoff(Units)

Ending Inventory

Stockouts (Units)

July

1300

1300

0

0

0

0

August

1150

1300

0

0

150

0

September

1100

1150

0

150

200

0

October

1600

1100

0

50

0

300

November

1900

1600

500

0

0

300

December

1900

1900

300

0

0

0

Total

800

200

350

600

Stockout happened in October as Inventory+production was not sufficient to meet demand so, stockout of (1600-1100-200) = 300 units happened

Total hiring cost = 800*35 = 28000
Total lay-off cost = 200*60 = 12000

Total inventory carrying cost = 350* 30 = 10500